OpenStudy (anonymous):

In your last 23 basketball games, you attempted 101 free throws and made 66. Find the experimental probability that you make a free throw. Write the probability as a percent, to the nearest tenth of a percent. A. 65.3% B. 69.8% C. 69.7% D. 65.7%

jigglypuff314 (jigglypuff314):

of the total 101 throws, 66 were made part/whole = 66/101 =

OpenStudy (anonymous):

A

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Ms. Wong is redecorating her office. She has a choice of 7 colors of paint, 4 kinds of curtains, 3 colors of carpet, and 2 styles of furniture. How many different ways are there to redecorate if she chooses two different colors of paint, one kind of curtain, one color of carpet, and one style of furniture? A. 168 ways B. 1,008 ways C. 1,176 ways D. 23 ways

jigglypuff314 (jigglypuff314):

7*7*4*3*2=

OpenStudy (anonymous):

C

jigglypuff314 (jigglypuff314):

yes :)

jigglypuff314 (jigglypuff314):

how many results were there total?

OpenStudy (anonymous):

i think its B

jigglypuff314 (jigglypuff314):

yes :)

jigglypuff314 (jigglypuff314):

how many odd numbers are there between 1 and 10?

OpenStudy (anonymous):

5

jigglypuff314 (jigglypuff314):

so each person has a 5 out of 10 chance of choosing an odd number. but the chance of both of them both choosing an odd number is (5/10)*(5/10)

OpenStudy (anonymous):

C

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

The probability that Shania is on time for school is 1/2. Find the probability that Shania arrives on time for school for the next 5 days. Express your answer as a percent, to the nearest tenth of a percent. A. 50% B. 15.6% C. 11.4% D. 3.1%

jigglypuff314 (jigglypuff314):

multiply the probability by itself the number of times it is continued, so (1/2)(1/2)(1/2)(1/2)(1/2)=

OpenStudy (anonymous):

D

jigglypuff314 (jigglypuff314):

yes :)

jigglypuff314 (jigglypuff314):

how many options are there per lottery number?

OpenStudy (anonymous):

8

jigglypuff314 (jigglypuff314):

... how many numbers are there 0 to 9

OpenStudy (anonymous):

C

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

yay :)

OpenStudy (anonymous):

Out of 1,100 discs tested, 13 are defective. Estimate the number of defective discs in a batch of 41,000. A. about 485 discs B. about 970 discs C. about 4,850 discs D. about 49 discs

jigglypuff314 (jigglypuff314):

\[\frac{ 13 }{ 1100 }= \frac{ x }{ 41000 }\]

OpenStudy (anonymous):

A

jigglypuff314 (jigglypuff314):

yes :)

jigglypuff314 (jigglypuff314):

sorry, but I am not sure how to help you with this one :/

OpenStudy (anonymous):

okay :/ its fine, but thank you

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@Loser66 @Mertsj